OpenStudy (anonymous):

1.13 × 10^1 + 9.1 × 10^-1

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

10^1=10 10^-1=0.1

OpenStudy (anonymous):

so really the equation is 1.13 x 10 + 9.1 x 0.1 ?

OpenStudy (anonymous):

so that means that this 4.1 x 10^4 - 1 x 10^3 is really 4.1 x 40 - 1 x 30 ?

OpenStudy (unklerhaukus):

10^4=10000≠40

OpenStudy (anonymous):

10^x stands for 10 power x. so 10^4 is 10 000 as noted by @UnkleRhaukus.

OpenStudy (anonymous):

so 10^4 does = 40 ?

OpenStudy (anonymous):

Some of this is coming back to me, I just want to be sure.

OpenStudy (anonymous):

no. Basically the x is the number of zeros after 1. 10^x is 1 and then x zeros.

OpenStudy (unklerhaukus):

\[10^4=10\times10\times10\times10=100\times100=10 000\]

OpenStudy (anonymous):

ohh okay, thank you (: